OpenStudy (anonymous):

A mug full to the brim with coffee weighs 520 grams. When it is half full, it weighs 400 grams. How many grams of coffee will it hold?

OpenStudy (anonymous):

let mug weight=a and full coffee=b a+b=520 a+b/2=400 subtracting b/2=120 b=240. full coffee =240 gm